Progress with Disk cylinder tooling Shielding foil added to cylinder Thermal feedthrough prepared and fitted Cylinder and stub wings assembled onto tooling (July) 4 disk alignment tools made by Manchester Disk grabber-design by Nikhef-manufactured by Sheffield/Lancaster Disk fixations from Nikhef and Liverpool-delivered alignment target holders designed (Nikhef) and manufactured (Lancaster) Disk insertion, alignment (X,Y), capture+removal (August) alignment better than 100 m Wheels+jacks added and tooling+cylinder have been moved in and out of the cold room (Sept)
Cylinder inside the cold room Moving into the cold room
Status of tooling-Oct 2 nd disk alignment, capture and removal has been performed with D7 populated with services –alignment checked with dummy masses –used interface to disk mini-cylinder –grounding: mini-cylinder and frame grounding –FSI spool+ ground foil handling –work with scaffolding around cylinder Work to be done –Thermal cycle of tooling+cylinder in the cold room –Removal and insertion of the beam to be demonstrated –grabber needs to be improved-stronger springs
Mounting disk onto cylinder from mini-cylinder Mounted relative to cylinder using pins FSI spool sitting on mini-cylinder
move FSI spool back onto PPF0 using PPF0 location screws Remove FSI spool and place outside patch panel Ground foils
Service status LMTs Testing of UK tapes at Glasgow, first 15 tested Delivery ~100/week, Taiwan to finish production 10/12 TIs Will be tested as part of LMT-TI harness at Glasgow String test LV/HV test cable TPPC TI LMT PPF0 wiggly tape hybrid/module at Liverpool 2 adjacent channels on pre-production TPPC tested First 16 TPPCs to be delivered to Liveerpool this week/next week for D9 Fibres Lancaster has volunteered to take on reception testing Fibres for disk testing (D1) at Liverpool, D9 delivered and tested ok, length issue DCS for D9 will be delivered by RB this week Cooling pipes Assume no cooling pipes initially and connect direct to disk as for disk testing Retro-fit cooling pipes at Liverpool/CERN when components ready and connect at STFT for testing Cooling-replicate disk testing system and connect direct to the disk-or similar Parts delivered and assembly underway at Liverpool
TPPC string test String test Test cable TPPC TI LMT harness PPF0 wiggly tape hybrid direct to hybrid String test Cooling HEX same design as for disk testing
Service handling Thermal plug delivered and fitted (Sheffield) PPF1 trays for LMTs ready (RAL) PPF1 tray holders (Sheffield) Delivered, being sent for anodising Other parts for LMT handling in progress at (Nikhef+Lancaster) Temporary PPF1 for fibres Lancaster have volunteered to make this based on prototype being made at Nikhef At Liverpool we need to ensure we can connect in the cold roomlooks good, measurement of fibre length indicated that fibres reach close to window in the thermal plug
Cooling connection captive ! Helicoflex not O-ring LMT connection checked connection with pcb by measuring thermistor use clip on light and dental mirror to check connections problem with some mis-alignments of guides passed back to disk makers, they will check alignment with pcbs can be fixed in-situ but to be avoided Strain relief at aperture
bendy fibres?! cable clamps (D3-D9 on cylinder) fibre clamps (D3-D9 on cylinder)
Temporary patch panel (TPPC) and TI connections
Testing during assembly Ensure connections from PPF0 on disk to TPP on thermal plug are well made before moving into cold roomlimited access! –LMT connections: check temperature through LV, checks connection at PPF0 –Check fibres and LMTs by sending trigger burst using SCT DAQ (RODDAQ being used to test disk) Aim to do module by module in parallel with service laying
Schedule-ambitious! 10/11 Assemble service handling –awaiting final parts Week of 15/11+22/11 leak check and test cooling, move into cold room and test with D7 –cooling being assembled –waiting for DCS cables (due this week) End November ready to install D9 and mount services
Issues Cooling and test of disk with services in cold room –cooling being set up for testing LMT delivery and making harnesses –delivery of D9 tapes by end this week ? –tape testing at harness preparation at Glasgow Problem with infininion connector at PPF0 –cannot be fixed in the cylinder Experience !
Testing in cold room Based on experience with disks Move into cold room and cable (1d) Check cabling (1.5d) Cooling down (1d) Confirmation tests (5d) Characterisation (5d) Warm-up (with I-V) (1d) Total ~15d per disks Can we run cold overnight ?!